### v2

# Bayesian KalmanNet: Quantifying Uncertainty in Deep Learning Augmented Kalman Filter

Abstract
Recent years have witnessed a growing interest in tracking algorithms that augment kf with dnn. By transforming kf into trainable deep learning models, one can learn from data to reliably track a latent state in complex and partially known dynamics. However, unlike classic kf, conventional dnn-based systems do not naturally provide an uncertainty measure, such as error covariance, alongside their estimates, which is crucial in various applications that rely on kf-type tracking. This work bridges this gap by studying error covariance extraction in dnn-aided kf. We begin by characterizing how uncertainty can be extracted from existing dnn-aided algorithms and distinguishing between approaches by their ability to associate internal features with meaningful kf quantities, such as the kg and prior covariance. We then identify that uncertainty extraction from existing architectures necessitates additional domain knowledge not required for state estimation. Based on this insight, we propose Bayesian KalmanNet, a novel dnn-aided kf that integrates Bayesian deep learning techniques with the recently proposed KalmanNet and transforms the kf into a stochastic machine learning architecture. This architecture employs sampling techniques to predict error covariance reliably without requiring additional domain knowledge, while retaining KalmanNet’s ability to accurately track in partially known dynamics. Our numerical study demonstrates that Bayesian KalmanNet provides accurate and reliable tracking in various scenarios representing partially known dynamic systems.
I Introduction
Tracking the latent state of a dynamic system is one of the most common task in signal processing [2]. Often, one is required not only to track the state, but also to provide a reliable assessment of the uncertainty in its estimate. For instance, navigation systems must not only provide the instantaneous position but also characterize its accuracy [3]. Uncertainty quantification is inherently provided by classic algorithms, such as the kf and the ekf (ekf), that rely on the ability to faithfully describe the dynamics as a closed-form ss (ss) model. Such model-based algorithms track not only the state, but also compute the error covariance [4].
The emergence of deep learning gave rise to a growing interest in using dnn to carry out tasks involving time sequences [5]. Various dnn architectures can learn from data to track dynamic systems without relying on full knowledge of an underlying ss model. These include generic rnn [6, Ch. 10] and attention mechanisms [7], as well as rnn-type architectures inspired by ss models [8] and kf processing [9]. Such dnn are trained end-to-end without using any statistical modeling.
When one has partial knowledge of the underlying dynamics, it can be leveraged by hybrid model-based/data-driven designs [10, 11, 12]. Existing approaches to exploit partial domain knowledge can be roughly divided following the machine learning paradigms of generative and discriminative learning [13]. Generative approaches are ss model-oriented, using dnn to learn the missing aspects of the underlying ss model, i.e., for system identification, while adopting parameterizations varying from physics-based ss models [14, 15] to fully dnn-based ones [16, 17, 18]. The learned models are then used by classical algorithms, such as the ekf, thus typically preserving the limitations of these methods to, e.g., Gaussian temporally independent noises [19].
The discriminative strategy is task-oriented, learning to directly output the state estimate, following the conventional usage of dnn for classification and regression [6]. Existing designs vary in the knowledge they require on the underlying model. For settings with relatively high level of domain knowledge, e.g., when one can approximately characterize the dynamics as a linear Gaussian ss model, dnn can be enhance classic kf by providing correction terms [20] or pre-processing [21] trained alongside the tracking algorithm. In scenarios with only partially known ss models, dnn augmented kf were proposed, converting the kf into a machine learning algorithm [22, 23, 24, 25, 26, 27, 28, 29]. Doing so bypasses the need to estimate the missing aspects of the ss model, learning to track via end-to-end training. In particular, the KalmanNet architecture [22] and its variants [23, 24, 26, 25] substitute the kg (kg) computation of the ekf with a dnn. The resulting algorithm was shown to enable accurate tracking in partially known and complex dynamics, while facilitating combination with pre-processing [30] and downstream state post-processing [31, 32, 33]. However, as such methods learn to estimate the state, they do not naturally provide the error covariance, for which there is rarely "ground-truth".
Uncertainty extraction is typically quite challenging in generic deep learning [34]. Even for classification dnn, that are designed to output a probability mass function, learned probabilities are often not faithful for uncertainty extraction [35], while in estimation tasks, generic dnn do not quantify uncertainty. For dnn-aided kf, several specific extensions were suggested to provide error covariance. For instance, the rkn (rkn) proposed in [9] was trained to output an error estimate in addition to the state estimate, while [36] showed that when applying KalmanNet of [22] in linear ss models with full column rank observation matrix and Gaussian measurement noise, the error covariance can be extracted from internal features of the architecture. These specific studies motivate a unifying study on error covariance extraction in dnn-aided kf.
In this work we study discriminative dnn-aided tracking with uncertainty quantification for both linear and non-linear partially known ss models. We extensively examine the extraction of error covariance from existing architectures. We categorize dnn that can learn to track in partially known ss model into model-based deep learning designs, where one can relate internal features to the kg or prior covariance, and black-box architectures, where only the output has an operational meaning. We show how the desired error covariance can be extracted in each setting: for black-box dnn, error covariance estimation necessitates providing additional output neurons; for architectures whose internal features can be attributed to the prior covariance and the kg, we identify how and in which conditions can its internal features be mapped into error covariance prediction. We present two training methods which encourage such architectures to support error covariance estimation.
Our analysis identifies that reliable extraction of the error covariance from existing dnn-aided algorithms is feasible only in some settings, e.g., when tracking fewer state variables than observed ones. Moreover, doing so requires additional knowledge of the ss model that is not required for state tracking. To overcome this, we propose a novel dnn-aided tracking algorithm that supports uncertainty quantification termed Bayesian KalmanNet. Our proposed algorithm draws inspiration from Bayesian deep learning [37, 38, 39], deviating from conventional frequentist learning to augment an ekf with a stochastic dnn to compute the kg. Doing so enables tracking the with the same level of partial ss modeling needed by conventional KalmanNet and its variants, while providing reliable uncertainty measures by sampling multiple dnn realizations.
We compare Bayesian KalmanNet and our extension of existing dnn-aided kf for uncertainty extraction in a qualitative and a quantitative manner. For the latter we consider both a synthetic and a physically compliant navigation setting. Our results show that Bayesian KalmanNet consistently provides reliable tracking of the state along with faithful error covariance prediction under different levels of accuracy in knowledge of the dynamics, and that incorporating domain knowledge in model-based deep learning designs systematically yields the most reliable estimate of the error covariance.
The rest of this paper is organized as follows: Section II briefly reviews ss models and the ekf. Extensions of existing dnn-aided kf for uncertainty extraction are presented in Section III. Bayesian KalmanNet is derived in Section IV, Section V details our numerical study, and Section VI provides concluding remarks.
[…]
II System Model and Preliminaries
In this section we review the system model and recall some preliminaries. We first present the ss model in Subsection II-A and discuss how it is used for model-based and dnn-based tracking in Subsections II-B-II-C, respectively. We formulate the problem of uncertainty-aware tracking in Subsection II-D, and review basics in Bayesian deep learning in Subsection II-E.
II-A SS Model
We consider a dynamic system described as a non-linear ss model in discrete time. This model describes the relationship between a latent state vector $\mathbf{x}_{t}\in\mathbb{R}^{m}$ and observation vector $\mathbf{y}_{t}\in\mathbb{R}^{n}$ for each time instance $t\in\mathbb{N}$. In a non-linear ss model, this relationship is generally represented as
$\mathbf{x}_{t}$ $=\mathbf{f}(\mathbf{x}_{t-1})+\boldsymbol{w}_{t},$ (1a)
$\mathbf{y}_{t}$ $=\mathbf{h}(\mathbf{x}_{t})+\boldsymbol{v}_{t}.$ (1b)
In (1), $\mathbf{f}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is the state evolution function; $\mathbf{h}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is the observation function; and $\boldsymbol{w}_{t},\boldsymbol{v}_{t}$ are mutually and temporally independent zero-mean noises with covariances $\mathbf{Q}$ and $\mathbf{R}$, respectively.
While various tasks are associated with ss models, including, e.g., smoothing and imputation, we focus on tracking. The tracking (filtering) task refers to estimating the state $\mathbf{x}_{t}$ from current and past observations $\{\mathbf{y}_{\tau}\}_{\tau\leq t}$. Performance is measured via the mse (mse) between the estimated state and the true one. Namely, for a sequence of length $T$, the time-averaged mse of an estimate $\{\hat{\mathbf{x}}_{t}\}_{t=1}^{T}$ is
${\rm MSE}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left\{\left\|\hat{\mathbf{x}}_{t}-{\mathbf{x}}_{t}\right\|^{2}\right\}.$ (2)
II-B Model-Based Tracking
The characterization of dynamic systems via the ss model (1) gives rise to various algorithms for tracking the state. Specifically, given a full and accurate description of the functions $\mathbf{f}(\cdot)$, $\mathbf{h}(\cdot)$, and characterization of the distribution of the state and measurement noises as being Gaussian with known covariances, a suitable tracking algorithm is the ekf [2, Ch. 10].
The ekf is comprised of prediction and measurement update steps. On each time step $t$, it starts by propagating the first-order statistical moments of the state and observations via
$\hat{\mathbf{x}}_{t|t-1}=\mathbf{f}(\hat{\mathbf{x}}_{t-1|t-1}),\quad\hat{\mathbf{y}}_{t|t-1}=\mathbf{h}(\hat{\mathbf{x}}_{t|t-1}).$ (3)
Letting $\hat{\mathbf{F}}_{t}$ and $\hat{\mathbf{H}}_{t}$ be the Jacobians (partial derivatives) of $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$, evaluated at $\hat{\mathbf{x}}_{t-1|t-1}$ and $\hat{\mathbf{x}}_{t|t-1}$, respectively, the prior covariances are set to
$\boldsymbol{\Sigma}_{t|t-1}$ $=\hat{\mathbf{F}}_{t}\boldsymbol{\Sigma}_{t-1|t-1}\hat{\mathbf{F}}_{t}^{\top}+\mathbf{Q},$ (4a)
$\mathbf{S}_{t|t-1}$ $=\hat{\mathbf{H}}_{t}\boldsymbol{\Sigma}_{t|t-1}\hat{\mathbf{H}}_{t}^{\top}+\mathbf{R}.\vspace{-0.1cm}$ (4b)
The observations are fused with the predictions as
$\hat{\mathbf{x}}_{t|t}=\hat{\mathbf{x}}_{t|t-1}+\mathbf{K}_{t}({\mathbf{y}}_{t}-\hat{\mathbf{y}}_{t|t-1}),$ (5)
where $\mathbf{K}_{t}$ is the kg given by
$\mathbf{K}_{t}=\boldsymbol{\Sigma}_{t|t-1}\hat{\mathbf{H}}_{t}^{\top}\mathbf{S}^{-1}_{t|t-1}.$ (6)
The error covariance is then computed as
$\boldsymbol{\Sigma}_{t|t}$ $=(\mathbf{I}-\mathbf{K}_{t}\hat{\mathbf{H}}_{t})\boldsymbol{\Sigma}_{t|t-1}.$ (7)
The ekf copes with the non-linearity of the ss model by linearizing the state evolution and measurement functions when propagating the second-order moments in (4). Alternative approaches employ different approximations for propagating these moments in non-linear ss models, including the unscented transform (utilized for that purpose by the ukf (ukf)  [40]), as well as the cubature [41] and Gauss-Hermite deterministic or stochastic quadrature rules [42, 43]. While these model-based algorithms may not necessarily be mse optimal (as opposed to linear Gaussian ss models), they $(i)$ inherently provide an estimate of both the state via (5) and the estimation uncertainty via (7); and $(ii)$ rely on accurate description of the dynamics as a known ss model as in (1).
II-C DNN-Aided Tracking
[…]
dnn, and particularly architectures suitable of time sequences such as rnn, can learn to track in dynamics where the ss model is complex and unknown [5]. dnn used for tracking take as input the current observation $\mathbf{y}_{t}$ at each time $t\in\mathbb{N}$. The output of such a dnn-aided filter with trainable parameters ${\boldsymbol{\theta}}$ at time $t$, denoted $\hat{\mathbf{x}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\in\mathbb{R}^{m}$, is an estimate of the state $\mathbf{x}_{t}$, as illustrated in Fig. 1(A).
We focus on supervised settings, where learning is based on data comprised of pairs of state length $T$ state trajectories and observations, denoted
$\mathcal{D}=\big{\{}\{\mathbf{x}_{t}^{(i)},\mathbf{y}_{t}^{(i)}\}_{t=1}^{T}\big{\}}_{i=1}^{|\mathcal{D}|},$ (8)
where the superscript $(i)$ denotes $i$th data trajectory. In training, the dnn parameters ${\boldsymbol{\theta}}$ are tuned such that the state estimate best matches the data. Since performance is measured via the mse (2), a suitable loss function is the empirical $\ell_{2}$ loss
$\mathcal{L}_{\mathcal{D}}^{\ell_{2}}({\boldsymbol{\theta}})=\frac{1}{|\mathcal{D}|T}\sum_{i=1}^{|\mathcal{D}|}\sum_{t=1}^{T}\big{\|}\mathbf{x}^{(i)}_{t}-\hat{\mathbf{x}}_{t}(\mathbf{y}^{(i)}_{t},{\boldsymbol{\theta}})\big{\|}^{2}.$ (9)
While dnn can be trained using (9) to track $\mathbf{x}_{t}$ without requiring characterization of the ss model in (1), they do not provide an estimate of the error covariance, e.g., as the ekf computes $\boldsymbol{\Sigma}_{t|t}$ in (7). Existing approaches for explainability in dnn, typically involve additional data and conformal prediction tools for boosting verification within confidence regions [44], rather than providing instantaneous error covariance.
II-D Problem Formulation
We consider the problem of uncertainty-aware tracking, namely, the design of an algorithm that maps the observed $\mathbf{y}_{t}$ into estimates of both the state $\hat{\mathbf{x}}_{t}$ and its estimation uncertainty $\hat{\boldsymbol{\Sigma}}_{t}$. The former is evaluated via the mse in (2), while the latter is measured via the accuracy of the estimated variances, i.e.,
$\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}\left\{\left|\left[\hat{\boldsymbol{\Sigma}}_{t}\right]_{j,j}-{\rm Var}\left\{\left[{\mathbf{x}}_{t}\right]_{j}\right\}\right|\right\}.$ (10)
Model-based tracking algorithms estimate both the state and the error covariance when the dynamics are characterized by a known ss model. However, we focus on setting where the ss is only partially known, and one must cope with the following requirements:
  1. R1
    The state evolution and observation models, $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$, are approximations of the true dynamics, that are possibly mismatched.
  2. R2
    The distribution of the noises ${\boldsymbol{v}}_{t}$ and ${\boldsymbol{w}}_{t}$, which capture the inherent stochasticity in the dynamics, is unknown, and can be complex and non-Gaussian.
  3. R3
    As error covariance is a statistical quantity, it is unlikely to be provided in a dataset $\mathcal{D}$. Accordingly, we assume access to data that is labeled as in (8), namely, it is comprised of trajectories of observations and states but does not include the error covariance.
These requirements motivate exploring data-driven approaches for tackling tasks associated with ss models, allowing to both recover the state and the error covariance.
II-E Bayesian Deep Learning
Bayesian deep learning offers a framework to quantify uncertainty in dnn predictions. It treats the dnn parameters as random variables, and learns their posterior distribution given the data. This probabilistic approach provides a measure of uncertainty by sampling multiple realizations of the (random) dnn during inference [45].
In Bayesian deep learning, instead of learning a single set of parameters ${\boldsymbol{\theta}}$, we aim to learn the posterior distribution $p({\boldsymbol{\theta}}|\mathcal{D})$ over the parameters given the data $\mathcal{D}$. In the considered context of estimating a state $\mathbf{x}_{t}$ from observations $\mathbf{y}_{t}$, the state estimate $\hat{\mathbf{x}}_{t}$ and the associated uncertainty can then be obtained by marginalizing over this posterior distribution, i.e.,
[…]
To learn a stochastic dnn, one typically seeks a distribution on the weights, i.e., $p({\boldsymbol{\theta}}|\mathcal{D})$, that is dictated by some parameters ${\boldsymbol{\phi}}$. This can result in either an explicit formulation of the distribution, as in variational inference [46], or via alternative induced trainable stochasticity, as in Monte Carlo dropout [47]. Once dictated, the parameters of the dnn distribution, ${\boldsymbol{\phi}}$, are trained using the elbo (elbo) objective, which regularizes the expected loss with a term minimizing the kl (kl) divergence between the induced distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ and some pre-determined prior $p({\boldsymbol{\theta}})$, i.e.,
$\text{ELBO}_{\mathcal{D}}({\boldsymbol{\phi}})=\mathbb{E}_{q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\left[\log p(\mathcal{D}|{\boldsymbol{\theta}})\right]-\text{KL}(q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})|p({\boldsymbol{\theta}})).$ (12)
Once trained, the stochastic dnn is applied via ensembling method. Particularly, during inference $J$ realizations, denoted $\{{\boldsymbol{\theta}}_{j}\}_{j=1}^{J}$, are drawn in an i.i.d. manner from $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$. Then, inference is carried out by applying these $J$ dnn to the input, and combining their results. For instance, in our context of state tracking, the resulting estimate becomes
$\hat{\mathbf{x}}_{t}({\boldsymbol{y}}_{t},{\boldsymbol{\phi}})=\frac{1}{J}\sum_{j=1}^{J}\hat{\mathbf{x}}_{t}^{(j)}(\mathbf{y}_{t},{\boldsymbol{\theta}}_{j}),\quad\{{\boldsymbol{\theta}}_{j}\}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}q({\boldsymbol{\theta}}|{\boldsymbol{\phi}}).$ (13)
III Error Covariance Extraction in Frequentist DNN-Aided Tracking
The first part of our study focuses on how one can extract error covariance from existing dnn-based tracking algorithms, based on conventional frequentist training (i.e., learning of deterministic dnn). We divide existing architectures by their interpretable features that can be related to error covariance in Subsection III-A. Then, we propose dedicated training objectives to encourage these architectures to allow extraction of faithful uncertainty measures in Subsection III-B, respectively, and provide a discussion in Subsection III-C.
III-A Error Covariance Indicative Features
Different dnn-aided algorithms provide different features that can be used to extract the error covariance. We divide dnn-aided systems accordingly, considering black-box methods, where only the outputs are observed, as well as model-based deep learning algorithms following the ekf operation, where one can possibly relate internal features to the prior covariance and kg.
III-A1 Output Features
In conventional black-box dnn, one can only assign an interpretable meaning to the input and output of the system. In such cases, in order to estimate the error covariance, additional outputs are required (see Fig. 1(B)); these outputs should represent the estimated covariance denoted $\hat{\boldsymbol{\Sigma}}_{t}$. This can be achieved by having the dnn output (in addition to $\hat{\mathbf{x}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$), a vector $\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\in\mathbb{R}_{+}^{m}$, as done in the rkn method [9], setting the estimated error covariance to be
$\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})={\rm diag}(\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})).$ (14)
III-A2 KG and Prior Covariance Features
Several hybrid model-based deep learning learn to track in partially known ss models by augmenting the ekf with dnn. Some algorithms, such as skn (skn) [23], learn to compute the kg, denoted $\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$, by treating an internal feature as an estimate of the prior covariance, denoted $\hat{\boldsymbol{\Sigma}}_{t|t-1}(\mathbf{y}_{t},{\boldsymbol{\theta}})$. Following the ekf (4)–(7), these features can be combined to estimate the error covariance as illustrated in Fig. 1(C) via
$\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})=(\mathbf{I}-\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\hat{\mathbf{H}}_{t})\hat{\boldsymbol{\Sigma}}_{t|t-1}(\mathbf{y}_{t},{\boldsymbol{\theta}}).$ (15)
The formulation in (15) is based on the update step of the ekf in (7), which implicitly assumes that the kg is the mse optimal linear estimator. As there is no guarantee that this indeed holds for a learned dnn-aided kg, one can still estimate the error covariance via the Joseph form
$\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$ $=\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\mathbf{R}\hat{\mathbf{K}}_{t}^{\top}(\mathbf{y}_{t},{\boldsymbol{\theta}})+(\mathbf{I}-\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\hat{\mathbf{H}}_{t})$
×$\times\hat{\boldsymbol{\Sigma}}_{t|t-1}(\mathbf{y}_{t},{\boldsymbol{\theta}})(\mathbf{I}-\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\hat{\mathbf{H}}_{t})^{\top}.$ (16)
III-A3 KG Features
The formulations in (15)-(16) rely on the ability to extract both the prior covariance and the kg as internal features. dnn-aided systems such as KalmanNet [22] learn to compute the kg from data without recovering the prior covariance. There, if $\tilde{\mathbf{H}}_{t}\triangleq\big{(}\hat{\mathbf{H}}_{t}^{\top}\hat{\mathbf{H}}_{t}\big{)}^{-1}$ exists, then, by combining (4b)-(6), the prior covariance can be recovered via
$\!\hat{\boldsymbol{\Sigma}}_{t|t\!-\!1}(\mathbf{y}_{t},{\boldsymbol{\theta}})\!=\!(\mathbf{I}\!-\!\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\hat{\mathbf{H}}_{t})^{-1}\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})\mathbf{R}\hat{\mathbf{H}}_{t}\tilde{\mathbf{H}}_{t}.$ (17)
Using the recovered prior covariance in (17), the desired error covariance is obtained by either (15) or (16). This approach is illustrated in Fig. 1(D).
Remark 1.
Converting the kg into the prior covariance in (17) requires $\tilde{\mathbf{H}}_{t}$ to exist, i.e., $\hat{\mathbf{H}}_{t}$ has full column rank. Yet, even when some of the columns of $\hat{\mathbf{H}}_{t}$ are zero or $n<m$, error covariance matrix elements with respect to the same state variables can still be recovered by considering the sub-matrices of $\hat{\mathbf{H}}_{t}$, $\hat{\boldsymbol{\Sigma}}_{t|t-1}$, and $\hat{\boldsymbol{\Sigma}}_{t}$ corresponding to these variables. For instance, in navigation scenarios with noisy position measurements, one can still extract the position error covariance while also tracking velocity.
III-B Training to Predict the Error Covariance
The indicative features detailed in Subsection III-A allow dnn-aided systems to provide $\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$, viewed as an estimate of the error covariance. However, training via (9), which solely observes the state estimate, does not explicitly encourage $\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$ to accurately match the error covariance. Accordingly, we consider two alternative loss measures that overcome the absence of a ground truth covariance, one that is based on evaluating the empirically averaged error variance, and one that utilizes a Gaussian prior of the estimated state.
Figure 1: Considered tracking approaches: $(A)$ standard black-box dnn; $(B)$ dnn with error covariance output (e.g., [9]); $(C)$ dnn-aided kf, estimated posterior covariance and kg (e.g., [23]); and $(D)$ dnn-aided kf, learned kg (e.g., [22]).
III-B1 Empirical Error Variance
A straight-forward approach to account for the error covariance in the loss is to introduce an additional term that compares $\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$ to the instantaneous empirical estimate. Specifically, by defining the error vector
$\boldsymbol{e}_{t}(\mathbf{y}_{t},\mathbf{x}_{t},{\boldsymbol{\theta}})\triangleq\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}}),$ (18)
the loss includes an additive term that compares the the predicted error with its empirical valuate obtained from (18). By defining the second moment loss over the whole dataset as:
$\mathcal{L}_{\mathcal{D}}^{\rm M2}({\boldsymbol{\theta}})=\sum_{i=1}^{|\mathcal{D}|}\sum_{t=1}^{T}\sum_{j=1}^{m}\left|[\boldsymbol{e}_{t}(\mathbf{y}^{(i)}_{t},\mathbf{x}^{(i)}_{t},{\boldsymbol{\theta}})]_{j}^{2}-[\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})]_{j,j}\right|.$ (19)
The resulting empirical loss used for training is
$\mathcal{L}_{\mathcal{D}}^{\rm Emp}({\boldsymbol{\theta}})=(1-\beta)\cdot\mathcal{L}_{\mathcal{D}}^{\ell_{2}}({\boldsymbol{\theta}})+\beta\cdot\mathcal{L}_{\mathcal{D}}^{\rm M2}({\boldsymbol{\theta}}).$ (20)
where $\beta\in[0,1]$ balances the contribution of each term. The hyperparameters $\beta$ can be gradually increased as training progresses to first learn to track and then also encourage reliable error prediction. Specifically, when setting $\beta=0$, then training with (20) coincides with training with the standard $\ell_{2}$ loss.
III-B2 Gaussian Prior
An alternative loss that accounts for error covariance without explicitly comparing predicted and empirical values leverages Gaussian priors. Particularly, one can train to both track and provide error covariance by imposing a Gaussian prior on the tracked state, as done in [9]. Under this assumption, a suitable loss is the log-likelihood of a Gaussian distribution with mean $\hat{\mathbf{x}}_{t}(\mathbf{y}^{(i)}_{t},{\boldsymbol{\theta}})$ and covariance $\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})$, i.e.,
$\mathcal{L}_{\mathcal{D}}^{\rm GP}({\boldsymbol{\theta}})=\frac{1}{|\mathcal{D}|T}\sum_{i=1}^{|\mathcal{D}|}\sum_{t=1}^{T}$ $\boldsymbol{e}_{t}^{\top}(\mathbf{y}^{(i)}_{t},\mathbf{x}^{(i)}_{t},{\boldsymbol{\theta}})\hat{\boldsymbol{\Sigma}}_{t}^{-1}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})$
×$\times\boldsymbol{e}_{t}(\mathbf{y}^{(i)}_{t},\mathbf{x}^{(i)}_{t},{\boldsymbol{\theta}})$
$+\log\det\left(\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})\right).$ (21)
We note that when the error covariance $\hat{\boldsymbol{\Sigma}}_{t}$ is not estimated, but is instead assumed to be a scaled identity, then training with the loss in (III-B2) is equivalent to training with the standard $\ell_{2}$ loss.
III-C Discussion
We identified three different approaches to extract error covariance in discriminative dnn-aided tracking in non-linear ss models. These approaches vary based on the architecture employed and the level of interpretability it provides. It is emphasized though that for all considered approaches, one must still alter the learning procedure for the extracted features to be considered as uncertainty measures. To encourage learning to accurately predict the error, we propose two training objectives. This examination of error covariance extraction in dnn-aided kf combines new approaches, while encompassing techniques used for specific schemes (e.g., (III-B2) used by rkn [9]), and extending previous findings (e.g., (17) extends [36] to non-linear models).
The main distinction between the approaches lies in their domain knowledge, i.e., which parts of the ss model are needed to track and provide error covariance, as summarized in Table I. Specifically, black-box dnn do not require any model knowledge. Nonetheless, as empirically demonstrated in Section V, using black-box architectures with additional outputs regarded as the error covariance often leads to inaccurate tracking and unreliable error prediction, even when trained with a covariance-oriented loss as the ones proposed in Subsection III-B. Model-based deep architectures that learn the kg require only approximate knowledge of $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$ to track [23, 22]. However, to extract covariance, $\mathbf{h}(\cdot)$ should be accurately known (as, e.g., (15) relies on $\hat{\mathbf{H}}_{t}$), where $\mathbf{R}$ is also needed when the prior covariance is not tracked, as in KalmanNet. Accordingly, requirements R1 and R2 are not fully met by these methods. Moreover, the latter also requires the linearized observation model (i.e., $\tilde{\mathbf{H}}_{t}$) to have full column rank. When these requirements are met, it is empirically shown in Section V that, when training via the losses proposed in Subsection III-B, one can both accurately track the latent state and reliably predict its covariance from the internal features of the dnn-aided algorithm.
IV Bayesian KalmanNet
As detailed in the previous section, and empirically demonstrated in Section V, model-based deep learning architectures based on the KalmanNet algorithm can simultaneously provide accurate state tracking alongside reliable error prediction. Nonetheless, achieving the latter requires deviating from how these dnn-aided tracking algorithms are designed (e.g., in [22] and [23]) in the following aspects:
  1. A1
    Additional domain knowledge that is not required merely for state tracking is needed for covariance extraction (e.g., full characterization of the observations model (1b)), and the family of dynamic systems where one can extract uncertainty is limited (e.g., $\tilde{\mathbf{H}}_{t}$ has to be non-singular).
  2. A2
    The training objective should be altered to encourage reliable uncertainty extraction.
In this section we propose a dnn-aided learning algorithm that preserves the ability of model-based deep learning algorithms such as KalmanNet to accurately track, in partially known ss models, while providing uncertainty measures without being subject to the limitations in A1, thus fully meeting requirements R1-R3. We build upon the observation in A2, namely, that error covariance extraction already necessitates altering the learning objective. Instead of just altering the learning objective, we modify the learning paradigm, deviating from conventional frequentist learning into Bayesian learning of stochastic dnn. Our proposed algorithm, coined Bayesian KalmanNet, is detailed in Subsection IV-A, with its training procedure formulated in Subsection IV-B, while Subsection IV-C provides a discussion.
IV-A Bayesian KalmanNet Tracking Algorithm
As the name suggests, Bayesian KalmanNet is based on the KalmanNet algorithm of [22], while employing Bayesian dnn. Accordingly, the algorithm uses a stochastic dnn to compute the kg of the ekf. The usage of a Bayesian dnn results in Monte Carlo sampling of $J$ different KalmanNet realizations during inference, that are used to recover both the state and its predicted error via ensemble averaging.
Architecture: We employ a Bayesian dnn with parameters ${\boldsymbol{\theta}}$ whose distribution is parameterized by ${\boldsymbol{\phi}}$. Each dnn realization maps the indicative features $\Delta\mathbf{x}_{t-1}\triangleq\hat{\mathbf{x}}_{t-1}-\hat{\mathbf{x}}_{t-2}$, $\Delta\mathbf{y}_{t}\triangleq\mathbf{y}_{t}-\mathbf{y}_{t|t-1}$ and $\Delta\tilde{\mathbf{y}}_{t}\triangleq\mathbf{y}_{t}-\mathbf{y}_{t-1}$ into the kg $\hat{\mathbf{K}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}})$.
In our numerical study, we employ Architecture 1 of [22], comprised of gru (gru) cells with input and output fc (fc) layers. Stochasticity is achieved by integrating dropout into the fc layers [47], with $\boldsymbol{p}$ denoting the Bernoulli distribution parameter assigned with each neuron of these layers. The trainable parameters dictating the dnn distribution, denoted ${\boldsymbol{\theta}}$, thus encompass a setting of the dnn parameters, denoted ${\boldsymbol{\theta}}_{0}$, as well as the dropout neuron-wise probability $\boldsymbol{p}$, i.e., ${\boldsymbol{\phi}}=[{\boldsymbol{\theta}}_{0},\boldsymbol{p}]$. In each realization, the dnn is applied with weights ${\boldsymbol{\theta}}_{0}$ with the fc layers employing a dropout mask randomized with per-element distribution $\boldsymbol{p}$.
Uncertainty-Aware Tracking: The proposed augmentation of the ekf with Bayesian dnn allows to track as in KalmanNet, while leveraging ensembling to predict the estimation error. When tracking is launched, we sample $J$ i.i.d. realizations of the kg computations dnn from $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$, denoted $\{{\boldsymbol{\theta}}_{j}\}_{j=1}^{J}$, with $J$ being a hyperparameter. For instance, when employing Monte Carlo dropout, each dnn realization is obtained from the initial fixed weight ${\boldsymbol{\theta}}_{j}$ and a realization of the dropout masks sampled with multivariate Bernoulli distribution $\boldsymbol{p}$.
On each time instance $t$, the first-order moments of the state and the observation are predicted based on the (possibly approximated) knowledge of $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$ via (3). These estimated moments are used to form the input features to each of the $J$ dnn realizations, obtaining $J$ kg, which are in turn used to produce $J$ separate state estimates, denoted $\{\hat{\mathbf{x}}_{t}^{j}\}_{j=1}^{J}$. The mean and variance of these estimates provide the state estimate and its associated uncertainty, respectively, via
$\hat{\mathbf{x}}_{t}$ $=\frac{1}{J}\sum_{j=1}^{J}\hat{\mathbf{x}}_{t}^{(j)}(\mathbf{y}_{t},{\boldsymbol{\theta}}_{j}),$ (22a)
$\hat{\boldsymbol{\Sigma}}_{t}$ $=\frac{1}{J}\sum_{j=1}^{J}(\hat{\mathbf{x}}_{t}^{(j)}-\hat{\mathbf{x}}_{t})(\hat{\mathbf{x}}_{t}^{(j)}-\hat{\mathbf{x}}_{t})^{\top}.$ (22b)
The resulting operation is summarized as Algorithm 1 (in which, for brevity, we write the output of the $j$th kg dnn realization at time $t$ as $\hat{\mathbf{K}}^{j}_{t}$), and illustrated in Fig. 2.
Figure 2: Bayesian KalmanNet - Architecture
Init : dnn distribution ${\boldsymbol{\phi}}$; ensemble size $J$
Input : Initial state $\mathbf{x}_{0}$
1 Set $\hat{\mathbf{x}}_{0}^{j}\equiv\mathbf{x}_{0}$;
2
3for $t=1,2,\ldots$ do
4      Receive new observation $\mathbf{y}_{t}$;
5
6      Sample $\{{\boldsymbol{\theta}}_{j}\}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$;
7
8      for $j=1,2,\ldots,J$ do
9            Predict $\hat{\mathbf{x}}_{t|t-1}^{j}$ and $\hat{\mathbf{y}}_{t|t-1}^{j}$ from $\hat{\mathbf{x}}_{t-1}$ via (3);
10
11            Compute kg $\hat{\mathbf{K}}^{j}_{t}$ via dnn ${\boldsymbol{\theta}}_{j}$;
12
13            Update $\hat{\mathbf{x}}_{t}^{j}=\hat{\mathbf{x}}_{t|t-1}^{j}+\hat{\mathbf{K}}^{j}_{t}({\mathbf{y}}_{t}-\hat{\mathbf{y}}_{t|t-1}^{j})$
14      Compute state estimate $\hat{\mathbf{x}}_{t}$ via (22a);
15
16      Predict estimation error $\hat{\boldsymbol{\Sigma}}_{t}$ via (22b);
17
Algorithm 1 Bayesian KalmanNet
IV-B Training Bayesian KalmanNet
The trainable parameters of Bayesian KalmanNet are the distribution parameters ${\boldsymbol{\phi}}$. In principle, one can train Bayesian KalmanNet as a standard Bayesian regression dnn, evaluating it based on the augmented ekf output using the elbo objective in (12). Nonetheless, as we are particularly interested in accurate error covariance recovery (which is not explicitly accounted for in the elbo formulation), we next propose a dedicated training loss, which is also geared towards our usage of Monte Carlo dropout for inducing stochasticity.
We follow the specialization of the elbo objective to concrete Monte Carlo based dnn proposed in [47], while altering the data matching term to incorporate explicit encouraging of reliable error variance recovery, as proposed in Subsection III-B for frequentist learning. To formulate this, let $L$ denote the number of layers in the dnn in which dropout is incorporated, and let $\{\lambda_{l}\}_{l=1}^{L}$ denote their indices. In the implementation detailed in Subsection IV-A, these indices correspond to the fc layer of KalmanNet’s Architecture 1. We write the parameters of layer $\lambda$ of ${\boldsymbol{\theta}}$ as ${\boldsymbol{\theta}}^{\lambda}$ and its dropout coefficient as $p^{(\lambda)}$. The proposed loss function, based on the labeled data in (8), is formulated as
$\mathcal{L}_{\mathcal{D}}^{\rm BKN}({\boldsymbol{\phi}}=[{\boldsymbol{\theta}}_{0},\boldsymbol{p}])=$
$\mathbb{E}_{{\boldsymbol{\theta}}\sim q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\{\mathcal{L}_{\mathcal{D}}^{\rm Emp}({\boldsymbol{\theta}})\}+\sum_{l=1}^{L}\mathcal{L}_{\lambda_{l}}^{\rm KL}({\boldsymbol{\theta}}_{0}^{\lambda_{l}},p^{(\lambda_{l})}),$ (23)
where $\mathcal{L}_{\lambda_{l}}^{\rm DO}({\boldsymbol{\phi}})$ computes the kl-based regularization term of the layer $\lambda_{l}$ based on the learned multivariate Bernoulli distribution, with pre-determined prior in (12) being a maximal entropy Bernoulli distribution. The layer-wise loss term is given by [47]
$\mathcal{L}_{\lambda}^{\rm KL}({\boldsymbol{\theta}},p)=c_{1}\cdot\frac{\|{\boldsymbol{\theta}}\|^{2}}{1-p}-c_{2}\cdot\mathcal{H}(p)\cdot n^{(\lambda)}.$ (24)
In (24), $n^{(\lambda)}$ is the number of input features to the $\lambda$th layer; $c_{1}$ and $c_{2}$ are constant coefficients (hyperparameters); and $\mathcal{H}(p)$ is the entropy of a Bernoulli random variable with probability $p$, defined as
$\mathcal{H}(p)=-p\log(p)-(1-p)\log(1-p).$ (25)
This loss in (23) guides the tuning of ${\boldsymbol{\phi}}$ when training Bayesian KalmanNet to both extract an accurate state estimation and error covariance (by taking the expectation of our proposed empirical error variance as the data matching term), while encouraging randomness in the prior on the dropout probability $\boldsymbol{p}$.
TABLE I: Domain knowledge requirements comparison
Method Feature $\mathbf{F}(\cdot)$ $\mathbf{H}(\cdot)$ $\mathbf{R}$ $\mathbf{Q}$ Output features e.g., RKN State None None None None Error cov None None None None Bayesian KalmanNet State Partial Partial None None Error cov Partial Partial None None Learn KG and Prior e.g., skn State Partial Partial None None Error cov Partial Full None None Learn KG e.g., KalmanNet State Partial Partial None None Error cov Partial Full Full None Model-based e.g., EKF State Full Full Full Full Error cov Full Full Full Full
IV-C Discussion
Bayesian KalmanNet is designed to allow dnn-aided tracking in partially known ss models with error prediction. The extensions to existing frequentist KalmanNet-type methods proposed in Section III, which map the interpretable internal features into the error covariance, follow the mathematical steps used by the ekf in such computations, and thus necessitate additional domain knowledge not required by KalmanNet for state tracking. Bayesian KalmanNet bypasses the need for such excessive knowledge by adopting ss model-agnostic Bayesian deep learning techniques for error covariance extraction, combined with the ss model-aware KalmanNet for state tracking. This results in Bayesian KalmanNet requiring the same level of partial ss model knowledge as needed by KalmanNet and skn for state estimation, while also providing error prediction, thus meeting R1-R3. As numerically reported in Section V, Bayesian KalmanNet achieves reliable state tracking and error prediction in various settings with different levels of domain knowledge.
The incorporation of Bayesian deep learning techniques can be viewed as trading computational complexity for uncertainty measures. This induced complexity also translates into additional inference latency, as reported in Section V. Specifically, the application of Bayesian KalmanNet can be viewed as $J$ realizations of KalmanNet with different parameterization, thus increasing the computational burden by a factor of $J$ at each time instance. In that sense, Bayesian KalmanNet bears similarity to alternative tracking algorithm that implement multiple Kalman-type filters in parallel, e.g., [48], though the conventional approach uses a different ss model with deterministic parameters for each filter, while Bayesian KalmanNet uses multiple stochastic realizations of the filter obtained for the same (partially known) ss model, following the task-oriented discriminative learning paradigm [13].
Our design of Bayesian KalmanNet gives rise to numerous possible avenues for extension. While our formulation accommodates a generic Bayesian dnn, our implementation and training objective focus on the usage of Monte Carlo dropout. Our formulation differs than standard Bayesian dnn as all realizations are based on the same prior. One can possibly consider alternative Bayesian parameterizations, based on, e.g., variational inference with explicit priors [46], as well as its incorporation with different KalmanNet-type architectures, e.g., [23, 24, 26, 25]. Moreover, the combination of Bayesian deep learning and KalmanNet indicates that such methodologies can also be employed in alternative forms of model-based deep learning [12]. We leave these extensions for future work.
V Numerical Evaluation
In this section we numerically compare the above techniques to extract error covariance in dnn-aided kf. We first detail the considered algorithms implemented and our utilized performance measures in Subsection V-A. Then, we detail a set of experiments dedicated to evaluating state and covariance estimation in different ss models in Subsections V-B-V-D. We conclude by comparing inference time in Subsection V-E.
V-A Experimental Setup
V-A1 Tracking Algorithms
We henceforth compare the following tracking algorithms [^1] The source code and all hyperparameters used can be found online at https://github.com/yonatandn/Uncertainty-extraction-in-Model-Based-DL:
  • rkn, using the code provided in [9], representing dnn with dedicated output features.
  • skn, using the code provided in [23], representing model-based deep learning with learned kg and prior covariance.
  • KalmanNet, using both Architecture 1 (KNetV1) and Architecture 2 (KNetV2) of [22].
  • Bayesian KalmanNet (termed BKN for brevity), based on KNetV1 with dropout incorporated into its fc layers, taking $J=20$ realizations.
  • The fully model-based ekf.
Unless stated otherwise, we use (15) to convert the kg and prior covariance into error covariance. We train rkn and skn with (III-B2); KalmanNet is trained with (9); while Bayesian KalmanNet is trained using (23). All data-driven algorithms are trained with the same data comprised of $|\mathcal{D}|=300$ trajectories.
V-A2 Data generation
Unless stated otherwise, throughout the follow experimental study, we set the observation noise covariance $\mathbf{R}$ to take the form $r^{2}\mathbf{I}$, with $1/r^{2}=n/{\rm Tr}(\mathbf{R})$ denoting the snr (snr). The ratio between the state evolution and the observation noise variances is fixed to $0.01$, such that changes in the snr affect both the state evolution and the observation noises. To perform some statistical tests and evaluation, mc simulation was performed, with $N=100$ trajectories per scenario.
V-A3 Performance Measures
The tracking algorithms are evaluated in terms of the following performance measures: $(i)$ The empirical estimation error, i.e., state estimate mse (2); and $(ii)$ The predicted error covariance, namely, ${\rm Tr}(\hat{\boldsymbol{\Sigma}}_{t})$. Performance is thus measured not only in state estimation error, but also on the ability to predict this error from the recovered covariance.
To assess the credibility of the estimators, we also evaluate the anees (anees) measure [49, 3, 50, 51]. This measure combines state estimation and error covariance prediction in a single measure. for an estimator outputting state estimates and covariances $\{\hat{\mathbf{x}}_{t},\hat{\boldsymbol{\Sigma}}_{t}\}_{t=1}^{T}$, the anees is given by
${\rm ANEES}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left\{(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\hat{\boldsymbol{\Sigma}}_{t}^{-1}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})\right\},$ (26)
where the stochastic expectation is computed by empirical averaging over the test data. The closer to the anees is to being $1$, the more credible the estimator, with values surpassing $1$ representing overconfidence, and lower values corresponding to conservative predictions.
Additional metrics used to evaluate the different methods at a given time instance $t$ over a test set $\mathcal{D}$ are the apec (apec) and the eec (eec). The apec is given by
[…]
while the eec is given by
[…]
We say that the algorithm is a better uncertainty estimator as its uncertainty prediction (apec along time) is closer to to the true error covariance (eec along time).
V-A4 Mismatch Setup
We consider the following forms of mismatches in the model knowledge:
  • Process noise mismatch: a dataset generated with a process noise covariance matrix $\mathbf{Q}_{\rm data}=100\cdot\mathbf{Q}_{\rm model}$ where $\mathbf{Q}_{\rm model}$ is the process noise covariance matrix known to the algorithm.
  • Measurement noise mismatch: a dataset generated with a measurement noise covariance matrix $\mathbf{R}_{\rm data}=100\cdot\mathbf{R}_{\rm model}$ where $\mathbf{R}_{\rm model}$ is the measurement noise covariance matrix known to the algorithm.
  • Model mismatch: in the cv (cv) ss, data was generated using a ca (ca) model, while testing the different methods with a cv ss.
V-B Linear Canonical SS Model
We first simulate a synthetic two-dimensional linear ss model with Gaussian process and measurement noises governed by a controllable canonical state transition matrix, defined in [22, Sec. IV.B]. Here, the state transition matrix ${\mathbf{F}}$ and measurement matrix ${\mathbf{H}}$ are
${\mathbf{F}}=\frac{1}{\sigma_{\text{max}}}\begin{bmatrix}1&1\\ 0&1\end{bmatrix},\quad{\mathbf{H}}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$
where $\sigma_{\text{max}}$ is the largest singular value of $\hat{\mathbf{F}}_{t}$ ensuring stability.
[…]
When introducing a mismatch in the measurement noise covariance $\mathbf{R}$ and evaluating the filter performance using the anees metric, as shown in Fig. 3(d), it is evident that bkn achieves the most favorable results. bkn provides a credible error covariance estimation, meaning it maintained reliability. The ekf and KalmanNet produce comparable results, indicating that both methods offer similar levels of credibility in their error covariance predictions. In contrast, the rkn struggles to provide credible estimates, leading to unreliable uncertainty predictions.
Turning to the mse results, illustrated in Fig. 3(c), it is clear that the skn outperforms other methods by achieving the lowest estimation error across various measurement noise covariance levels. This indicates that skn consistently provided the most accurate state estimates, regardless of the noise magnitude. The rkn, however, only performed well at higher $r^{2}$ values, suggesting that it struggles in high snr where accurate uncertainty modeling is more critical. All KalmanNet based methods, including bkn, outperform the ekf.
The results in this ss model emphasize that for a variety of snr, even for a linear state transition and measurement functions, both the frequentist and the Bayesian methods combining model knowledge and deep learning achieve better results than both model agnostic deep learning and model based without deep learning.
((a)) mse for mismatched $\mathbf{Q}$, all the model aware methods managed to track the state with a very similar mse. the model agnostic rkn has failed to maintaine the same performance
((b)) anees for mismatched $\mathbf{Q}$, Both the frequentist and bayesian KNet have managed to mainatain a stable error covariance (uncertainty) estimation throughout the different snr.
((c)) mse for mismatched $\mathbf{R}$, the skn performance was superior while both the frequentist and bayesian model-based deep learning methods managed to track the state with a very similar mse.
((d)) anees for mismatched $\mathbf{R}$, the bkn consistently provided the most accurate and stable uncertainty estimates, whereas the model agnostic rkn’s prediction was the mostly inaccurate.
Figure 3: Canonical linear ss model - mismatched process noise $\mathbf{Q}$ and measurement noise $\mathbf{R}$ covariances
V-C Pendulum SS Model
Next, we consider a pendulum with mass $m=1$ and length $\ell=1$. The pendulum exhibits nonlinear behavior due to the gravitational force $g$ acting on it. We describe the system using the angle $\theta$ and angular velocity $\dot{\theta}$ as the state variables, sampled at interval of $\Delta t$. Let the state vector $\mathbf{x}_{t}$ be defined as:
$\mathbf{x}_{t}=\begin{bmatrix}\theta_{t}\\ \dot{\theta}_{t}\end{bmatrix}.$
The dynamics of the pendulum can be modeled using Newton’s second law. For a pendulum subject to gravity, we can express this second-order equation as a system of first-order differential equations such that the discrete nonlinear state transition function $\mathbf{F}(\mathbf{x}_{t})$ can be written as:
$\mathbf{f}(\mathbf{x}_{t})=\begin{bmatrix}\theta_{t}+\dot{\theta}_{t}\Delta t\\ \dot{\theta}_{t}-\frac{g}{\ell}\sin(\theta_{t})\Delta t\end{bmatrix}.$
Assuming that $\mathbf{x}_{t}$ at each time step is small enough to allow a local linearization, the Jacobian matrix of $\mathbf{F}(\mathbf{x}_{t})$ is
$\hat{\mathbf{F}}_{t}=\begin{bmatrix}1&\Delta t\\ -\frac{g}{\ell}\cos(\theta_{t})\Delta t&1\end{bmatrix}.$
We assume the pendulum’s position in Cartesian coordinates, such that the observation matrix ${\mathbf{h}}(\mathbf{x}_{t})$ for converting state to measurement is given by:
${\mathbf{h}}(\mathbf{x}_{t})=\begin{bmatrix}\ell\cos(\theta_{t})\\ \ell\sin(\theta_{t})\end{bmatrix}.$
Here, $\hat{\mathbf{H}}(\mathbf{x}_{t})$ is non-full columns rank. This mean that the frequentist model-based methods are unable to extract an error covariance matrix as explained in Remark 1, ans thus are not tested throughout this scenario.
The achieved performance measures are reported in Fig. 4. When examining the mse in Fig. 4(a), it can be seen that the ekf fails to produce an accurate state estimation compared to all other methods, especially at higher snr. The anees results in Fig. 4(b) highlight that beside the skn, all algorithms have a stable credibility with a minor overconfidence throughout different snr, whereas the skn shows some instability in its credibility. The rkn provides the most accurate error covariance estimates compared to the other methods when introducing a mismatch in the process noise covariance $\mathbf{Q}$.
Under mismatched $\mathbf{R}$, the mse results in Fig. 4(c) show that in terms of estimation error the ekf performs the worst, as its MSE remains significantly higher than the other methods across the entire range of $1/r^{2}$. This indicates that the ekf struggles to maintain accurate state estimates, particularly in the presence of higher measurement noise. Its performance gap widens considerably as the noise mismatch increases, showing that the ekf is less robust under such conditions. The dnn-aided bkn, skn, and rkn all perform much better, achieving comparable results for most of the range.
[…]
((a)) mse for mismatched $\mathbf{Q}$, All the deep learning methods showed stable results but struggled slightly at moderate snr. The ekf, however, failed to maintain competitive performance.
((b)) anees for mismatched $\mathbf{Q}$, the different methods provided very similar results, beside the instability of the skn’s credibility for some snr.
((c)) mse for mismatched $\mathbf{R}$, All the deep learning methods showed stable results with a significantly better performance than the ekf in the sense of the estimated state accuracy.
((d)) anees for mismatched $\mathbf{R}$, the model based deep learning methods consistently provide the most accurate and stable uncertainty estimates compared to the model-free and model dependent methods.
Figure 4: Pendulum non-linear ss model - mismatched process noise $\mathbf{Q}$ and measurement noise $\mathbf{R}$ covariances
V-D Navigation SS Model
Our third scenario focuses on a more realistic ss as we track the location and velocity of a vehicle moving in a 2D plane. We assume its movement is in a cv manner. The system is modeled using the vehicle’s position and velocity in the $x$- and $y$-directions as state variables.
Let the state vector $\mathbf{x}_{t}$ and state transition matrix ${\mathbf{F}}$ be defined as:
$\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ y_{t}\\ \dot{x}_{t}\\ \dot{y}_{t}\end{bmatrix},\quad{\mathbf{F}}=\begin{bmatrix}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}$
where:
  • $x_{t}$ and $y_{t}$ represent the position of the vehicle in the $x$- and $y$-directions.
  • $\dot{x}_{t}$ and $\dot{y}_{t}$ represent the velocity of the vehicle in the $x$- and $y$-directions.
Assume we can observe the position of the vehicle in the $x$- and $y$-directions, but not its velocity. The observation matrix ${\mathbf{H}}$ maps the state vector to the measurements:
${\mathbf{H}}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}$
The results for this setting are reported in Fig. 5. When examining the mse in Fig. 5(a), it can be seen that both the ekf and the rkn not only fail to estimate the uncertainty, but also fail to produce an accurate state estimation compared to all other methods, especially at higher snr. The results in Fig. 5(b) shows that all methods resulted in an overconfident covariance prediction, although the both the frequentist and Bayesian model based deep learning methods manage to have a much better credibility with a more accurate uncertainty predictions.
((a)) mse for mismatched $\mathbf{Q}$, the model based deep learning methods consistently provide the most accurate state estimates compared to the model-free and model dependent methods.
((b)) anees for mismatched $\mathbf{Q}$, Both the frequentist and bkn manage to maintain an accurate error covariance estimation throughout the different snr.
Figure 5: cv ss model - mismatched process noise $\mathbf{Q}$
In this section, we describe a single run of our constant velocity model under specific conditions. The true process noise covariance $\mathbf{Q}_{\rm data}$ was set such that the given process noise covariance $\mathbf{Q}_{\rm model}$ is underestimated by a factor of 100, i.e., $\mathbf{Q}_{\rm data}=100\cdot\mathbf{Q}_{\rm model}$. The noise covariance $\mathbf{R}$ was set to $10$, which is 100×100$\times$100 × greater than $\mathbf{Q}_{\rm model}$. This setup aims to analyze the system’s behavior under significant model-process mismatch.
Figure 6: x-axis estimate in a single run, cv ss model.
Figure 6 shows the predicted position x track of each method compared to the ground true trajectory. This figure is a qualitative measure that shows the ability to track the trajectory. As can be seen, the model agnostic rkn failed to track the state along time, whereas all the model aware methods provide a reasonable state estimation.
((a)) KalmanNet V1
((b)) Bayesian KalmanNet
((c)) KalmanNet V2
((d)) Split-KalmanNet
Figure 7: The eec and apec (as described in (27) and (28)) Vs. time for the model-based deep learning methods.
((a)) Extended Kalman Filter
((b)) Recurrent Kalman Network
Figure 8: The eec and apec (as described in 27 and 28) for the model-based ekf and model-agnostic rkn.
To compare the different methods’ performance in prediction the uncertainty, Figs. 7 and 8 show x-axis position apec compared to its eec. These metrics are achieved by performing a simulation with $N=100$ trajectories, and using (27) and (28) to calculate apec and eec respectively. As can be seen, the model-based deep learning methods achieve better results than the model-based EKF and the model-agnostic RKN.
The results under model mismatch due to sampling mismatch are reported in Fig. 9. Fig. 9(b) shows clearly that bkn has supplied the most reliable error covariance estimation when given a mismatched model. The model based deep learning frequentist methods also results in a reliable error covariance prediction, whereas the ekf fails to extract the uncertainty, outperforming only the model agnostic rkn. The results for the state estimation when given a mismatched model are very similar to the results in the error covariance estimation. Fig. 9(a) shows that the bkn has supplied the most reliable state estimation, the frequentist methods resulted in a reasonable state prediction, whereas the ekf and rkn failed to estimate an accurate state estimation.
((a)) mse for mismatched model, the model based deep learning methods consistently provide the most accurate state estimates compared to the model-free and model dependent methods.
((b)) anees for mismatched model, The frequentist KNets have managed to maintain an accurate error covariance estimations, although the bkn achieved the most credible error covariance estimation.
Figure 9: cv ss model - mismatched model
V-E Inference Time Comparison
The results presented so far evaluate the considered methods in terms of their accuracy and uncertainty estimation performance. As Bayesian deep learning induces excessive latency in inference compared with frequentist models due to its inherent ensembling, we next evaluate inference time. The results in Table II report the inference latency of the different filtering methods across both a linear (canonical) and non-linear (pendulum) scenario. The platform on which all algorithms were tested and runtimes were computed is an Intel(R) Core(TM) i7-10510U CPU @ 1.80GHz 2.30 GHz. In the canonical ss model, the ekf demonstrates the fastest performance of all model aware methods, with the lowest average inference time of $5.279$ milliseconds. The skn, KNet V1, and V2 methods exhibit comparable speeds, with marginal differences around $5.3$ milliseconds. These results suggest that these methods maintain efficient performance close to the traditional ekf, making them suitable for real-time applications where speed is critical.
The rkn achieves an average inference time of just $0.194$ milliseconds. This presumably stems from the efficiency of rnn software implementations in PyTorch. bkn requires substantially more time, with an inference time of $104.88$ milliseconds for the canonical model. This reflects the computational cost associated with more advanced uncertainty quantification and Bayesian learning techniques. While these methods enhance credibility and robustness, they introduce a significant computational burden, making them less ideal for time-sensitive tasks.
In the pendulum scenario, the inference times increase across all methods due to the model’s non-linearity. The ekf inference time rises to $13.842$ milliseconds, which is still faster than bkn at $112.461$ milliseconds. Notably, the KalmanNet versions do not have available inference times for the pendulum case, as the measurement matrix $\mathbf{H}$ in that ss is not full column rank, preventing the calculation of the error covariance. The skn retains its relatively low inference time of $13.906$ milliseconds, maintaining its appeal for applications requiring both computational efficiency and accuracy.
Overall, while the bkn offers substantial benefits in terms of credibility and robustness, these come at the cost of longer inference times. For real-time applications, a parallelization could be used to solve this issue, as the different inferences in the bkn are independent and thus could be computed in parallel.
TABLE II: Average Inference Time [msec]
Type Model EKF KNet V1 KNet V2 SKN RKN BKN Linear Canonical 5.279 5.303 5.279 5.315 0.194 104.88 Non-Linear Pendulum 13.842 N/A N/A 13.906 0.198 112.461
VI Conclusions
In this work, we studied uncertainty extraction in dnn-aided kf. We identified three techniques based on different indicative features, and proposed two learning schemes that encourage learning to predict the error. Moreover, we have suggested a new method termed bkn incorporating Bayesian deep learning to a dnn augmented kf. Our numerical results show that incorporating domain knowledge notably facilitates uncertainty extraction, and that bkn allows dnn-aided kf to both extract the state and predict the error accurately in non-linear possibly mismatched ss models.

### v3

@@
 I Introduction
-Tracking the latent state of a dynamic system is one of the most common task in signal processing [2].
+Tracking the latent state of a dynamic system is one of the most common tasks in signal processing [2].
 Often, one is required not only to track the state, but also to provide a reliable assessment of the uncertainty in its estimate.
 For instance, navigation systems must not only provide the instantaneous position but also characterize its accuracy [3].
+Moreover, uncertainty quantification in state estimation provides a measure of confidence in the predictions and enables informed, risk-aware decision-making, which is particularly critical in safety-sensitive domains.
 Uncertainty quantification is inherently provided by classic algorithms, such as the kf and the ekf (ekf), that rely on the ability to faithfully describe the dynamics as a closed-form ss (ss) model.
@@
 Existing designs vary in the knowledge they require on the underlying model.
-For settings with relatively high level of domain knowledge, e.g., when one can approximately characterize the dynamics as a linear Gaussian ss model, dnn can be enhance classic kf by providing correction terms [20] or pre-processing [21] trained alongside the tracking algorithm.
+For settings with relatively high level of domain knowledge, e.g., when one can approximately characterize the dynamics as a linear Gaussian ss model, dnn can enhance classic kf by providing correction terms [20] or pre-processing [21] trained alongside the tracking algorithm.
 In scenarios with only partially known ss models, dnn augmented kf were proposed, converting the kf into a machine learning algorithm [22, 23, 24, 25, 26, 27, 28, 29].
@@
 Our proposed algorithm draws inspiration from Bayesian deep learning [37, 38, 39], deviating from conventional frequentist learning to augment an ekf with a stochastic dnn to compute the kg.
-Doing so enables tracking the with the same level of partial ss modeling needed by conventional KalmanNet and its variants, while providing reliable uncertainty measures by sampling multiple dnn realizations.
+Doing so enables tracking the state with the same level of partial ss modeling needed by conventional KalmanNet and its variants, while providing reliable uncertainty measures by sampling multiple dnn realizations.
 We compare Bayesian KalmanNet and our extension of existing dnn-aided kf for uncertainty extraction in a qualitative and a quantitative manner.
@@
 II-E Bayesian Deep Learning
-Bayesian deep learning offers a framework to quantify uncertainty in dnn predictions.
+While the formulation of the ss model in Subsection II-A and the model-based tracking in Subsection II-B focuses on a Bayesian setting, from a deep learning perspective, the term Bayesian deep learning refers to a framework in which a stochastic modeling is employed on the dnn itself.
+Accordingly, we henceforth use the term frequentist within the context of dnn to refer to the standard deep learning approach in which the weights of a dnn are deterministic weights, whereas a Bayesian dnn models its weights as random variables governed by a learned posterior distribution [37].
+Specifically, Bayesian deep learning offers a framework to quantify uncertainty in dnn predictions.
 It treats the dnn parameters as random variables, and learns their posterior distribution given the data.
@@
 […]
-To learn a stochastic dnn, one typically seeks a distribution on the weights, i.e., $p({\boldsymbol{\theta}}|\mathcal{D})$, that is dictated by some parameters ${\boldsymbol{\phi}}$.
+To learn a stochastic dnn, one typically seeks to approximate a posterior distribution on the weights, i.e., $p({\boldsymbol{\theta}}|\mathcal{D})$, that is dictated by some parameters ${\boldsymbol{\phi}}$.
 This can result in either an explicit formulation of the distribution, as in variational inference [46], or via alternative induced trainable stochasticity, as in Monte Carlo dropout [47].
 Once dictated, the parameters of the dnn distribution, ${\boldsymbol{\phi}}$, are trained using the elbo (elbo) objective, which regularizes the expected loss with a term minimizing the kl (kl) divergence between the induced distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ and some pre-determined prior $p({\boldsymbol{\theta}})$, i.e.,
-$\text{ELBO}_{\mathcal{D}}({\boldsymbol{\phi}})=\mathbb{E}_{q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\left[\log p(\mathcal{D}|{\boldsymbol{\theta}})\right]-\text{KL}(q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})|p({\boldsymbol{\theta}})).$ (12)
+$\text{ELBO}_{\mathcal{D}}({\boldsymbol{\phi}})=\mathbb{E}_{q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\left[\log p({\boldsymbol{\theta}}|\mathcal{D})\right]-\text{KL}(q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})|p({\boldsymbol{\theta}})).$ (12)
 Once trained, the stochastic dnn is applied via ensembling method.
@@
 $\boldsymbol{e}_{t}(\mathbf{y}_{t},\mathbf{x}_{t},{\boldsymbol{\theta}})\triangleq\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}(\mathbf{y}_{t},{\boldsymbol{\theta}}),$ (18)
-the loss includes an additive term that compares the the predicted error with its empirical valuate obtained from (18). By defining the second moment loss over the whole dataset as:
+the loss includes an additive term that compares the predicted error with its empirical value obtained from (18). By defining the second moment loss over the whole dataset as:
 $\mathcal{L}_{\mathcal{D}}^{\rm M2}({\boldsymbol{\theta}})=\sum_{i=1}^{|\mathcal{D}|}\sum_{t=1}^{T}\sum_{j=1}^{m}\left|[\boldsymbol{e}_{t}(\mathbf{y}^{(i)}_{t},\mathbf{x}^{(i)}_{t},{\boldsymbol{\theta}})]_{j}^{2}-[\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})]_{j,j}\right|.$ (19)
@@
 $\mathcal{L}_{\mathcal{D}}^{\rm Emp}({\boldsymbol{\theta}})=(1-\beta)\cdot\mathcal{L}_{\mathcal{D}}^{\ell_{2}}({\boldsymbol{\theta}})+\beta\cdot\mathcal{L}_{\mathcal{D}}^{\rm M2}({\boldsymbol{\theta}}).$ (20)
-where $\beta\in[0,1]$ balances the contribution of each term.
-The hyperparameters $\beta$ can be gradually increased as training progresses to first learn to track and then also encourage reliable error prediction.
+The hyperparameter $\beta\in[0,1]$ in (20) is used to balance the trade-off between minimizing the tracking error and encouraging accurate prediction of the error covariance.
+In our numerical implementation detailed in Section V, $\beta$ is initialized with a small value at the beginning of training, so that the network initially focuses on learning to accurately track the state.
+As training progresses, $\beta$ is gradually increased, thereby placing more emphasis on the accuracy of the predicted uncertainty.
 Specifically, when setting $\beta=0$, then training with (20) coincides with training with the standard $\ell_{2}$ loss.
@@
 $+\log\det\left(\hat{\boldsymbol{\Sigma}}_{t}(\mathbf{y}_{t}^{(i)},{\boldsymbol{\theta}})\right).$ (21)
-We note that when the error covariance $\hat{\boldsymbol{\Sigma}}_{t}$ is not estimated, but is instead assumed to be a scaled identity, then training with the loss in (III-B2) is equivalent to training with the standard $\ell_{2}$ loss.
+We note that when the error covariance $\hat{\boldsymbol{\Sigma}}_{t}$ is not estimated, but is instead assumed to be a scaled identity, then training with the loss in (21) is equivalent to training with the standard $\ell_{2}$ loss.
 III-C Discussion
@@
 To encourage learning to accurately predict the error, we propose two training objectives.
-This examination of error covariance extraction in dnn-aided kf combines new approaches, while encompassing techniques used for specific schemes (e.g., (III-B2) used by rkn [9]), and extending previous findings (e.g., (17) extends [36] to non-linear models).
+This examination of error covariance extraction in dnn-aided kf combines new approaches, while encompassing techniques used for specific schemes (e.g., (21) used by rkn [9]), and extending previous findings (e.g., (17) extends [36] to non-linear models).
 The main distinction between the approaches lies in their domain knowledge, i.e., which parts of the ss model are needed to track and provide error covariance, as summarized in Table I.
@@
 realizations of the kg computations dnn from $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$, denoted $\{{\boldsymbol{\theta}}_{j}\}_{j=1}^{J}$, with $J$ being a hyperparameter.
-For instance, when employing Monte Carlo dropout, each dnn realization is obtained from the initial fixed weight ${\boldsymbol{\theta}}_{j}$ and a realization of the dropout masks sampled with multivariate Bernoulli distribution $\boldsymbol{p}$.
+The type of distribution represented by $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ is a design hyperparameter, as different families of distributions are considered in the Bayesian deep learning literature [46].
+In our numerical study reported in Section V we employ Monte Carlo dropout [47], where each dnn realization is obtained from the initial fixed weight ${\boldsymbol{\theta}}_{j}$ and a realization of the dropout masks sampled with multivariate Bernoulli distribution $\boldsymbol{p}$.
+However, the formulation of Bayesian KalmanNet is not restricted to Monte Carlo dropout, and can be combined with alternative distributions, e.g., Gaussian weights whose parameters are the first- and second-order moments.
 On each time instance $t$, the first-order moments of the state and the observation are predicted based on the (possibly approximated) knowledge of $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$ via (3).
@@
 $\hat{\boldsymbol{\Sigma}}_{t}$ $=\frac{1}{J}\sum_{j=1}^{J}(\hat{\mathbf{x}}_{t}^{(j)}-\hat{\mathbf{x}}_{t})(\hat{\mathbf{x}}_{t}^{(j)}-\hat{\mathbf{x}}_{t})^{\top}.$ (22b)
-The resulting operation is summarized as Algorithm 1 (in which, for brevity, we write the output of the $j$th kg dnn realization at time $t$ as $\hat{\mathbf{K}}^{j}_{t}$), and illustrated in Fig. 2.
+From a Bayesian deep learning perspective, the matrix $\hat{\boldsymbol{\Sigma}}_{t}$ in (22b) represents an empirical estimate of the predictive uncertainty, namely, the second-order moment of the posterior predictive distribution over the state $\mathbf{x}_{t}$ given the observations and training data [48].
+Specifically, each estimate $\hat{\mathbf{x}}_{t}^{(j)}$ in (22a) corresponds to a sample from a different model instantiation, i.e., from the distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ over the neural network weights.
+This ensemble of predictions approximates the posterior predictive distribution, and the sample covariance $\hat{\boldsymbol{\Sigma}}_{t}$ estimates the total predictive uncertainty.
+While there is no general guarantee that individual estimators $\hat{\mathbf{x}}_{t}^{(j)}$ are realizations of an unbiased estimator, the usage of its sample covariance for quantifying uncertainty is a standard and widely accepted approach in Bayesian deep learning.
+Accordingly, our proposed Bayesian KalmanNet uses $\hat{\boldsymbol{\Sigma}}_{t}$ in (22b) as an estimate of the error covariance.
+The resulting operation is summarized as Algorithm 1 (in which, for brevity, we write the output of the $j$th kg dnn realization at time $t$ as $\hat{\mathbf{K}}^{j}_{t}$), and illustrated in Fig.
+2.
 Figure 2: Bayesian KalmanNet - Architecture
@@
 We write the parameters of layer $\lambda$ of ${\boldsymbol{\theta}}$ as ${\boldsymbol{\theta}}^{\lambda}$ and its dropout coefficient as $p^{(\lambda)}$.
+Accordingly, the trainable parameters of Bayesian KalmanNet in such settings, denoted ${\boldsymbol{\phi}}$, include the weights of each of the layers ${\boldsymbol{\theta}}_{0}$, as well as the distribution parameters for the layers in which dropout is incorporated, denoted by $\boldsymbol{p}=[p^{\lambda_{1}},\ldots,p^{\lambda_{L}}]$.
 The proposed loss function, based on the labeled data in (8), is formulated as
@@
 $\mathbb{E}_{{\boldsymbol{\theta}}\sim q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\{\mathcal{L}_{\mathcal{D}}^{\rm Emp}({\boldsymbol{\theta}})\}+\sum_{l=1}^{L}\mathcal{L}_{\lambda_{l}}^{\rm KL}({\boldsymbol{\theta}}_{0}^{\lambda_{l}},p^{(\lambda_{l})}),$ (23)
-where $\mathcal{L}_{\lambda_{l}}^{\rm DO}({\boldsymbol{\phi}})$ computes the kl-based regularization term of the layer $\lambda_{l}$ based on the learned multivariate Bernoulli distribution, with pre-determined prior in (12) being a maximal entropy Bernoulli distribution.
-The layer-wise loss term is given by [47]
+where $\mathcal{L}_{\lambda_{l}}^{\rm KL}({\boldsymbol{\phi}})$ computes the kl-based regularization term of the layer $\lambda_{l}$ based on the learned multivariate Bernoulli distribution, with pre-determined prior in (12) being a maximal entropy Bernoulli distribution.
+The empirical loss term and the KL loss term in (23) are balanced through the non-negative hyperparameters $c_{1},c_{2}$, that appear in the layer-wise loss term, which is given by [47]
 $\mathcal{L}_{\lambda}^{\rm KL}({\boldsymbol{\theta}},p)=c_{1}\cdot\frac{\|{\boldsymbol{\theta}}\|^{2}}{1-p}-c_{2}\cdot\mathcal{H}(p)\cdot n^{(\lambda)}.$ (24)
-In (24), $n^{(\lambda)}$ is the number of input features to the $\lambda$th layer; $c_{1}$ and $c_{2}$ are constant coefficients (hyperparameters); and $\mathcal{H}(p)$ is the entropy of a Bernoulli random variable with probability $p$, defined as
+In (24), $n^{(\lambda)}$ is the number of input features to the $\lambda$th layer; and $\mathcal{H}(p)$ is the entropy of a Bernoulli random variable with probability $p$, defined as
 $\mathcal{H}(p)=-p\log(p)-(1-p)\log(1-p).$ (25)
 This loss in (23) guides the tuning of ${\boldsymbol{\phi}}$ when training Bayesian KalmanNet to both extract an accurate state estimation and error covariance (by taking the expectation of our proposed empirical error variance as the data matching term), while encouraging randomness in the prior on the dropout probability $\boldsymbol{p}$.
-TABLE I: Domain knowledge requirements comparison
+Having formulated the loss in (23), the training of Bayesian KalmanNet follows the conventional stochastic gradient based optimization of dnn, only with a Bayesian loss term.
+Accordingly, the overall training complexity is comparable to that of frequentist KalmanNet [22].
+TABLE I: Domain knowledge requirements comparison.
 Method Feature $\mathbf{F}(\cdot)$ $\mathbf{H}(\cdot)$ $\mathbf{R}$ $\mathbf{Q}$ Output features e.g., RKN State None None None None Error cov None None None None Bayesian KalmanNet State Partial Partial None None Error cov Partial Partial None None Learn KG and Prior e.g., skn State Partial Partial None None Error cov Partial Full None None Learn KG e.g., KalmanNet State Partial Partial None None Error cov Partial Full Full None Model-based e.g., EKF State Full Full Full Full Error cov Full Full Full Full
@@
 As numerically reported in Section V, Bayesian KalmanNet achieves reliable state tracking and error prediction in various settings with different levels of domain knowledge.
+The uncertainty measure extracted by Bayesian KalmanNet encapsulates the overall estimation error covariance.
+This uncertainty measure inherently comprises two components [48]: aleatoric uncertainty, resulting from the stochasticity of the system (e.g., process and measurement noises); and epistemic uncertainty, arising due to limited or mismatched knowledge within the estimator, particularly when the underlying dynamics are only partially known or when trained with a limited dataset.
+Classical model-based methods, such as the ekf, can reliably quantify aleatoric uncertainty when operating with an accurate ss model and well-characterized noise statistics, they underperform when the model is misspecified.
+Conversely, existing frequentist dnn-aided kf struggle to provide any form of uncertainty quantification, even with abundant training data.
+The approaches proposed for extending frequentist dnn-aided kf introduced in Section III are geared towards quantifying aleatoric uncertainty, aiming to recreate the covariance computations of the model-based ekf without explicitly propagating second-order moments.
+Accordingly, the resulting characterization, which does not encompass epistemic uncertainty, is expected to be sensitive to the amount of training data.
+Bayesian KalmanNet overcomes these limitations by providing uncertainty estimates that reflect both sources of uncertainty, thereby yielding more reliable and calibrated error covariances under data scarcity and model mismatch.
+This dual capability makes Bayesian KalmanNet particularly suitable for practical scenarios where both the dynamics and noise statistics are only partially known.
 The incorporation of Bayesian deep learning techniques can be viewed as trading computational complexity for uncertainty measures.
 This induced complexity also translates into additional inference latency, as reported in Section V.
-Specifically, the application of Bayesian KalmanNet can be viewed as $J$ realizations of KalmanNet with different parameterization, thus increasing the computational burden by a factor of $J$ at each time instance.
-In that sense, Bayesian KalmanNet bears similarity to alternative tracking algorithm that implement multiple Kalman-type filters in parallel, e.g., [48], though the conventional approach uses a different ss model with deterministic parameters for each filter, while Bayesian KalmanNet uses multiple stochastic realizations of the filter obtained for the same (partially known) ss model, following the task-oriented discriminative learning paradigm [13].
+The hyperparameter $J$ plays a key role in affecting both the performance, as increasing $J$ enhances the approximation of the posterior predictive distribution, while also increasing complexity.
+Specifically the application of Bayesian KalmanNet can be viewed as $J$ realizations of KalmanNet with different parameterizations, thus increasing the computational burden by a factor of $J$ at each time instance.
+The induced latency of Bayesian KalmanNet can be notably reduced using parallel computing, as each of the $J$ realizations can be applied simultaneously.
+We also note that this operation of Bayesian KalmanNet bears similarity to alternative tracking algorithm that implement multiple Kalman-type filters in parallel, e.g., [49], though the conventional approach uses a different ss model with deterministic parameters for each filter, while Bayesian KalmanNet uses multiple stochastic realizations of the filter obtained for the same (partially known) ss model, following the task-oriented discriminative learning paradigm [13].
 Our design of Bayesian KalmanNet gives rise to numerous possible avenues for extension.
-While our formulation accommodates a generic Bayesian dnn, our implementation and training objective focus on the usage of Monte Carlo dropout.
+Our design, as well as that considered for frequentist KalmanNet [22], assumes that the ss model in (1) does not change.
+Thus, our approach is not naturally transferable to time-varying dynamics, for which adaptive kf, often based on variational inference [50, 51, 52], are required.
+Previous works that considered the KalmanNet methodology in time-varying dynamics, which proposed adaptation based on hypernetworks [53] and unsupervised learning [54], focused on block-wise variations in the ss model.
+The extension of the KalmanNet methodology to time-varying dynamics is the area of ongoing research, which, inspired by recent advances in continual Bayesian learning [55, 56], can potentially benefit from the incorporation of Bayesian deep learning.
+Moreover, while our formulation accommodates a generic Bayesian dnn, our implementation detailed in Section V is based on an architecture proposed in [22], and training objective focus on the usage of Monte Carlo dropout.
 Our formulation differs than standard Bayesian dnn as all realizations are based on the same prior.
@@
 Unless stated otherwise, we use (15) to convert the kg and prior covariance into error covariance.
-We train rkn and skn with (III-B2); KalmanNet is trained with (9); while Bayesian KalmanNet is trained using (23).
-All data-driven algorithms are trained with the same data comprised of $|\mathcal{D}|=300$ trajectories.
+We train rkn and skn with (21); KalmanNet is trained with (9); while Bayesian KalmanNet is trained using (23).
+All data-driven algorithms are trained with the same data comprising of $|\mathcal{D}|=300$ trajectories.
+Training commences with the dnn weights initialized via standard Xavier initialization, and the Monte Carlo dropout parameters of BKN randomized uniformly in $[0.5,0.8]$, following the common practice in Bayesian dnn [47].
+The remaining learning hyperparameters were selected based on standard cross validation.
 V-A2 Data generation
-Unless stated otherwise, throughout the follow experimental study, we set the observation noise covariance $\mathbf{R}$ to take the form $r^{2}\mathbf{I}$, with $1/r^{2}=n/{\rm Tr}(\mathbf{R})$ denoting the snr (snr).
+Unless stated otherwise, throughout the following experimental study, we set the observation noise covariance $\mathbf{R}$ to take the form $r^{2}\mathbf{I}$, with $1/r^{2}=n/{\rm Tr}(\mathbf{R})$ denoting the snr (snr).
 The ratio between the state evolution and the observation noise variances is fixed to $0.01$, such that changes in the snr affect both the state evolution and the observation noises.
-To perform some statistical tests and evaluation, mc simulation was performed, with $N=100$ trajectories per scenario.
+All dnn-aided algorithms were trained on a range of snr rather than for each snr separately, to emphasize their robustness and ability to operate across different noise levels.
+To perform statistical tests and evaluation, mc simulation was performed, with $N=100$ trajectories per scenario.
 V-A3 Performance Measures
@@
 […]
-We say that the algorithm is a better uncertainty estimator as its uncertainty prediction (apec along time) is closer to to the true error covariance (eec along time).
+We say that the algorithm is a better uncertainty estimator as its uncertainty prediction (apec along time) is closer to the true error covariance (eec along time).
 V-A4 Mismatch Setup
@@
   • Model mismatch: in the cv (cv) ss, data was generated using a ca (ca) model, while testing the different methods with a cv ss.
+Our motivation for employing the considered mismatches, imposed on all algorithms that require knowledge of the ss model, is to evaluate sensitivity to incorrect modeling assumptions.
+Specifically, the noise mismatches (which affect just the ekf) serve to highlight the behavior of uncertainty quantification under severe mismatches, while the last model mismatch represents a realistic setup in which a target follows a more complicated state evolution model compared to the one used by the tracking algorithm.
 V-B Linear Canonical SS Model
@@
 In contrast, the rkn struggles to provide credible estimates, leading to unreliable uncertainty predictions.
+Comparing Figs.
+3(b) and 3(d) shows that frequentist methods can also achieve strong performance, particularly in scenarios where uncertainty can be reliably estimated.
+However, in cases where uncertainty estimation is inherently more challenging, e.g., where inaccuracies directly impact the observed data—frequentist methods relying on internal features may struggle.
+In such settings, Bayesian KalmanNet provides more robust and calibrated uncertainty estimates by explicitly modeling parameter uncertainty through ensemble predictions.
+Overall, both variants exhibit stable and credible uncertainty estimates across different SNR levels, and the observed differences are relatively minor.
 Turning to the mse results, illustrated in Fig.
-3(c), it is clear that the skn outperforms other methods by achieving the lowest estimation error across various measurement noise covariance levels.
+3(c), it is clear that the skn, which is particularly designed for settings with imbalances between state evolution and measurement noises [23], outperforms other methods by achieving the lowest estimation error across various measurement noise covariance levels.
 This indicates that skn consistently provided the most accurate state estimates, regardless of the noise magnitude.
-The rkn, however, only performed well at higher $r^{2}$ values, suggesting that it struggles in high snr where accurate uncertainty modeling is more critical.
+The rkn, however, only performed well at higher $r^{2}$ values, suggesting that it struggles in high snr with limited data sets, where accurate uncertainty modeling is more critical.
 All KalmanNet based methods, including bkn, outperform the ekf.
+Comparing the mse results in Fig.
+3(c) and the uncertainty results in Fig.
+3(d) indicates on the trade-off emerging by the need to balance both objectives during training.
+In such multi-objective settings, improving one aspect-—e.g., achieving lower MSE-—may slightly compromise the precision of uncertainty estimates, sometimes somewhat compromise the precision of uncertainty estimates, and in some cases hardly affect the latter.
 The results in this ss model emphasize that for a variety of snr, even for a linear state transition and measurement functions, both the frequentist and the Bayesian methods combining model knowledge and deep learning achieve better results than both model agnostic deep learning and model based without deep learning.
-((a)) mse for mismatched $\mathbf{Q}$, all the model aware methods managed to track the state with a very similar mse. the model agnostic rkn has failed to maintaine the same performance
+((a)) mse for mismatched $\mathbf{Q}$, all the model aware methods managed to track the state with a very similar mse. The model agnostic rkn has failed to maintain the same performance.
 ((b)) anees for mismatched $\mathbf{Q}$, Both the frequentist and Bayesian KNet have managed to mainatain a stable error covariance (uncertainty) estimation throughout the different snr.
@@
 The pendulum exhibits nonlinear behavior due to the gravitational force $g$ acting on it.
-We describe the system using the angle $\theta$ and angular velocity $\dot{\theta}$ as the state variables, sampled at interval of $\Delta t$.
+We describe the system using the angle $\theta$ and angular velocity $\dot{\theta}$ as the state variables, sampled at interval of $\Delta t=0.01$.
 Let the state vector $\mathbf{x}_{t}$ be defined as:
@@
 Here, $\hat{\mathbf{H}}(\mathbf{x}_{t})$ is non-full columns rank.
-This mean that the frequentist model-based methods are unable to extract an error covariance matrix as explained in Remark 1, ans thus are not tested throughout this scenario.
+This means that the frequentist model-based methods are unable to extract an error covariance matrix as explained in Remark 1, and thus are not tested throughout this scenario.
 The achieved performance measures are reported in Fig.
@@
 […]
+TABLE II: Performance metrics at snr $101/r^{2}=$ 10 [dB] with mismatched $\mathbf{R}$, Pendulum non-linear ss model.
+Metric EKF SKN RKN BKN
+MSE [dB] 23.8933 -7.5536 -7.7836 -9.2741
+log(ANEES) 4.1182 1.2452 0.2234 -0.0834
 ((a)) mse for mismatched $\mathbf{Q}$, All the deep learning methods showed stable results but struggled slightly at moderate snr. The ekf, however, failed to maintain competitive performance.
@@
 ${\mathbf{H}}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}$
+We generate a set of 100 trajectories, with 5 used for validation and 10 reserved for testing. The sampling interval parameter $\Delta t$ is normalized to unity.
 The results for this setting are reported in Fig.
@@
 Figure 6: x-axis estimate in a single run, cv ss model.
-Figure 6 shows the predicted position x track of each method compared to the ground true trajectory.
+Fig.
+6 shows the predicted position x track of each method compared to the ground truth trajectory.
 This figure is a qualitative measure that shows the ability to track the trajectory.
@@
 As Bayesian deep learning induces excessive latency in inference compared with frequentist models due to its inherent ensembling, we next evaluate inference time.
-The results in Table II report the inference latency of the different filtering methods across both a linear (canonical) and non-linear (pendulum) scenario.
+The results in Table III report the inference latency of the different filtering methods across both a linear (canonical) and non-linear (pendulum) scenario.
 The platform on which all algorithms were tested and runtimes were computed is an Intel(R) Core(TM) i7-10510U CPU @ 1.80GHz 2.30 GHz.
+For bkn (bkn), The latency values reflect the total execution time required to process all $J$ realizations sequentially, i.e., the actual inference latency experienced in a serial execution setting.
 In the canonical ss model, the ekf demonstrates the fastest performance of all model aware methods, with the lowest average inference time of $5.279$ milliseconds.
@@
 This reflects the computational cost associated with more advanced uncertainty quantification and Bayesian learning techniques.
-While these methods enhance credibility and robustness, they introduce a significant computational burden, making them less ideal for time-sensitive tasks.
+While these methods enhance credibility and robustness, they introduce a significant computational burden, making them less suited for time-sensitive tasks.
 In the pendulum scenario, the inference times increase across all methods due to the model’s non-linearity.
@@
 For real-time applications, a parallelization could be used to solve this issue, as the different inferences in the bkn are independent and thus could be computed in parallel.
-TABLE II: Average Inference Time [msec]
+While there is no consistent algorithm that achieves the best performance in both accuracy and uncertainty throughout the various settings considered in our numerical study, it is clearly demonstrated that incorporating domain knowledge notably facilitates uncertainty extraction.
+Moreover, it is systematically shown that bkn enables dnn-aided kf to both extract the state and predict the error accurately in nonlinear, possibly mismatched ss models, and that its reliability is maintained across all the considered settings.
+TABLE III: Average Inference Time [msec].
 Type Model EKF KNet V1 KNet V2 SKN RKN BKN Linear Canonical 5.279 5.303 5.279 5.315 0.194 104.88 Non-Linear Pendulum 13.842 N/A N/A 13.906 0.198 112.461